OpenStudy (anonymous):

HELP!!!!! let f(x)=lnsqrt((4x+6)/(9x+6)) find f'(x)

OpenStudy (zugzwang):

\[f(x)=\ln \sqrt{\frac{4x+6}{9x+6}}\] Is this it?

OpenStudy (anonymous):

Yes

OpenStudy (zugzwang):

Chain rule, do you know it?

OpenStudy (anonymous):

not very well in this problem I tried doing the quotient rule at first and epicaly failed

OpenStudy (zugzwang):

Ok, chain rule, as best as I can say it, first, try to determine the outermost function... In this case, it's ln, right?

OpenStudy (anonymous):

yes and the derivative of that is 1/x

OpenStudy (zugzwang):

That's right. So whatever is inside the ln, just treat it as one big x. So you have \[\frac{1}{\sqrt{\frac{4x+6}{9x+6}}}\]

OpenStudy (zugzwang):

We're not yet done, ok, this is just part of the process...

OpenStudy (anonymous):

So then would you apply the quotient rule to what is under the radical?

OpenStudy (zugzwang):

No, now that you've pretty much done away with the "ln" part, now, what is the outermost function? It's the square root, right? What's the derivative of a square root?

OpenStudy (anonymous):

1/2(4x+6/9x+6)^-1/2

OpenStudy (zugzwang):

That's right, and you multiply that to that other fraction we had a while ago. We're not yet done. Now what's the outermost function? It's the rational function, the (4x+6)/(9x+6), just get the derivative of that, and multiply it to the ones we already got, and you're done :)

OpenStudy (anonymous):

so for that inside we do use the quotient rule correct?

OpenStudy (zugzwang):

Yes, that's when you use the quotient rule.